# Perspective on Moreau–Yosida Regularization in Density-Functional Theory

### v1

@@ III.1 MY based on the universal density functional
The most straightforward way to introduce MY regularization into DFT follows from its definition in Eq. (15): By directly applying the method to the non-differentiable (but convex and lower semicontinuous) $F^{\lambda}(\rho)$, yielding the regularized universal density functional $F_{\varepsilon}^{\lambda}(x)$, which is well-defined and finite for all $x\in X$,
@@
where $\varepsilon>0$ is the smoothing parameter. An illustration of a regularized functional in one dimension together with the potentially non-convex $\tilde{F}^{\lambda}(x)$ and convex $F^{\lambda}(x)$ is shown in Figure 1. The regularized functional has a well-defined functional derivative (see v from the properties of MY regularization in Section II.3) and the regularization is lossless in the sense that it conserves the overall infimum of the functional (see ii). We next introduce the corresponding energy functional $E_{\varepsilon}^{\lambda}(v)$ as the convex conjugate of $F_{\varepsilon}^{\lambda}(x)$,
@@
Note that the resulting $E_{\varepsilon}^{\lambda}(v)$ is not the MY regularization of $E^{\lambda}(v)$, but instead the following relation holds (from Eq. (17), where the different sign is due to Eq. (32)),
@@
Since we subtract the strictly concave norm-square of $v$ from the concave $E^{\lambda}(v)$, the MY energy $E_{\varepsilon}^{\lambda}(v)$ is strictly concave. Conversely, it holds that the Legendre–Fenchel transform of a strictly concave energy functional is always differentiable. It is precisely this expression for the energy functional from where we start in the next motivation for MY regularization that is more physically informed.
With $v=-\nabla F_{\varepsilon}^{\lambda}(x)$ we now have a unique representing potential $v\in X^{*}$ for every $x\in X$. We have already written $\nabla F_{\varepsilon}^{\lambda}$ instead of the subdifferential $\partial F_{\varepsilon}^{\lambda}$ because we know that the regularized functional is differentiable. This then automatically yields a well-defined and unique mapping from ‘densities’ (elements of $X$) to potentials (elements of $X^{*}$) without any $v$-representability issue. We are not even limited to the $N$-representability set $D\subsetneq X$ any more, and Section III.3 will help to make sense of the generalized densities $x\in X$. Uniqueness of the potential also means we automatically have a Hohenberg–Kohn-type result.
In the Hilbert-space setting (Assumption 3), we obtain also a quantitative bound that shows how the difference between two potentials is bound by the difference between the corresponding densities, in form of the Lipschitz estimate from property ix,
@@
This bound is obviously lost in the limit $\varepsilon\to 0$, in agreement with the fact that, in standard DFT, two densities that lie very close to each other can come from vastly different potentials [51], making density–potential inversion notoriously difficult. A similar bounds is also possible in more general Banach spaces if the MY regularization is adapted to the ‘modulus of convexity’ of the space [52]. Furthermore, MY regularization allows for a well-defined formulation of the KS iteration procedure (Section V).
$\tilde{F}^{\lambda}(x)$$F^{\lambda}(x)$$F_{\varepsilon}^{\lambda}(x)$
@@ III.2 MY based on the total energy functional
A different viewpoint starts from the concave ground-state energy functional $E^{\lambda}(v)$ and introduces a strictly concave energy $E_{\varepsilon}^{\lambda}(v)$ by just subtracting $\varepsilon\phi(v)=\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$, i.e.,
@@
The convex conjugate of this strictly concave functional then yields the regularized density functional of Eq. (49) as shown by Eq. (17),
@@
We have thus defined MY regularization starting from the energy functional instead of the universal density functional. But talking about the total energy allows to give this manipulation a physical interpretation. Say, for any external potential $v\in X^{*}$, we define the energy functional
$E^{\mathrm{f}}(v)=-\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}.$ (55)
Note the minus sign that makes this functional concave, just like the concave $E^{\lambda}(v)$ of the matter system. Of course, to retrieve the correct physical electrostatic field energy, $\frac{\varepsilon_{0}}{2}\|\nabla v\|^{2}_{L^{2}}$, an appropriate choice of Banach space $X^{*}$ must be made; see Section VI.1. We then introduce the new energy functional (that could be interpreted as the total energy of matter and field)
@@
Regarding the negative sign in Eq. (55), remember that this is the (concave) energy functional and not the internal energy expression that is actually varied over. There, the additional term is $\tfrac{1}{2\varepsilon}\|x-\rho\|^{2}_{X}$ as seen in Eq. (49) and using the relation motivated in Eq. (57) below this is equal to $\tfrac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$, precisely the expression for the (positive) field energy. Viewed in this manner, it is the energy content of the external field that regularizes $F^{\lambda}(\rho)$. Conceptually, this is an extremely attractive viewpoint, since it explains the regularization as a physical effect.
@@ III.3 MY based on density–potential mixing
Remember that we defined the $N$-representability set $D\subsetneq X$ in Eq. (29) as the set of all $\rho\in X$ where $F_{\mathrm{DM}}^{\lambda}(\rho)<\infty$, known as the effective domain of $F_{\mathrm{DM}}^{\lambda}(\rho)$. For the sake of argument, we now suppose Assumption 4 under which $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$. Now even for $\rho\in D$, we can have issues with $\partial F^{\lambda}(\rho)$. First, $\partial F^{\lambda}(\rho)$ may be empty. In such cases, no potential $v\in X^{\ast}$ exists such that $v\in-\partial F^{\lambda}(\rho)\iff\rho\in\partial E^{\lambda}(v)$, implying that $\rho$ is not $v$-representable. Also, a small variation of $\rho$ can (and will) lead outside of the set of $v$-representable densities [35; 51] and there are examples on small lattice systems where even large variations of $v$ lead to almost (or even exactly) the same ground-state density $\rho$ [32; 53].
It would be better altogether if we could ignore this difficult notion of “$v$-representability” and instead give a physical interpretation to all elements $x\in X$. This is achieved with the inverse duality map $J^{-1}\colon X^{*}\to X$ as the canonical way of mapping potentials to densities. For any $v$-representable $\rho\in\partial E^{\lambda}(v)$ coming from $v\in X^{*}$ define the mixed density
@@
This definition is unique only if a Hohenberg–Kohn-type result is available so that the corresponding $v\in-\partial F^{\lambda}(\rho)$ is unique, but even without uniqueness we can make this assignment and will simply allow multiple such combinations. Choosing $\varepsilon>0$ small means that in order to get the generalized density, we only mix in a small quantity of the corresponding potential $v$. We note that the resulting $x$ does not need to be positive, normalizable, or have a finite kinetic energy, only the $v$-representable density $\rho$ has all those properties. With $v\in X^{*}$ as an external potential with assigned density $\varepsilon J^{-1}(v)$ that gets added to the internal electronic density of a quantum system, one could call $x$ a ‘bidensity’. It has been given different names before, such as ‘depot’ (for density–potential), ‘pseudodensity’ [1], or ‘quasidensity’ [54]. We will not commit to a singular nomenclature here, and instead refer to it as a general element $x\in X$ or a ‘mixed density’ when its decomposition like in Eq. (57) matters.
One physically motivated possibility is to choose the spaces such that $J$ corresponds to solving the Poisson equation of electrostatics and $\varepsilon J^{-1}(v)=-\varepsilon\Delta v$. This choice was previously suggested in the context of density–potential inversion [4]; it will be covered in more detail in Section VI.1 and has already found multiple use in ameliorating the Kohn–Sham method [55; 56; 57], albeit not with the connection to the topology of density and potential space.
@@ IV Density–potential inversion
The limit procedure in property vii of the MY regularization in Section II.3 that says $\lim_{\varepsilon\to 0}\nabla F_{\varepsilon}^{\lambda}(x)\in\partial F^{\lambda}(x)$ will be the basis for a method to calculate the corresponding external potential from a given ground-state density. The method itself is already known since a long time as the “ZMP method” [2], but our derivation as well as the accompanying interpretation is only feasible by resorting to MY regularization. We also add a more speculative subsection that makes an interesting and previously unexplored connection to the Dyson equation. In the section we will require Assumption 1 but sometimes, when it is mentioned, demand higher regularity of the involved spaces for additional results.
@@
whose stationarity condition $\partial E^{\lambda}(v)+\rho\ni 0$ gives $v\in-\partial F^{\lambda}(\rho)$ as an optimizer if $\rho$ is assumed $v$-representable. This approach to density–potential inversion was first applied to atoms by Colonna and Savin [58] in 1999 and by Teale, Coriani and Helgaker to molecules [59; 60; 61] in 2009. It sometimes called “Lieb optimization”, “Lieb maximization” [60; 61; 62] or the “dual problem” [63] in the literature. Convergence of this method can be facilitated by making the functional in the optimization strictly concave by subtracting $\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$, which corresponds to MY regularization,
@@
We use Eq. (58) to translate this to $\rho\in\partial E_{\varepsilon}^{\lambda}(v_{\varepsilon})$ which is equivalent to $v_{\varepsilon}=-\nabla F_{\varepsilon}^{\lambda}(\rho)$ by using the analogous relation to (39) for the regularized functionals. By property v from Section II.3, we can connect $v_{\varepsilon}$ to the proximal map,
@@ IV.1 Proximal-point iteration and relation to ZMP
As mentioned before, it is not guaranteed that this subdifferential is non-empty even for $\rho\in D\subsetneq X$ and, in non-standard settings for DFT (like DFT on finite lattices [32]), it may even be multi-valued. For the regularized functional, no such problems occur, and we can evaluate
$v_{\varepsilon}=-\nabla F_{\varepsilon}^{\lambda}(\rho),$ (65)
@@
Due to the possibility of degeneracy in the ground state, the next step $p_{i+1}$ is possible non-unique, which creates a possible issue that has not been explored up to date. If converged, this process yields the proximal point $p=\mathop{\mathrm{prox}}_{\varepsilon F^{\lambda}}(\rho)$ and we retrieve Eq. (63),
@@
where the resulting potential $v_{\varepsilon}$ depends on the choice of the smoothing parameter $\varepsilon$. Repeating this iterative scheme for a sequence of decreasing $\varepsilon>0$, we can extrapolate to the limit $\varepsilon=0$ and obtain $v_{\varepsilon}\rightharpoonup v$. Under the minimal Assumption 1, convergence $v_{\varepsilon}\rightharpoonup v$ is weak; under the stronger Assumption 2, convergence is strong. By virtue of vii, the resulting potential $v$ is the element of $\partial F^{\lambda}(\rho)$ with the smallest norm.
The benefit of this method is that we do not need to evaluate the (usually unknown) $F^{\lambda}(\rho)$, relying instead on $\partial E^{\lambda}(v)$ by solving the Schrödinger equation. This density-potential inversion procedure using MY regularization was previously detailed in Penz, Csirik, and Laestadius [4].
@@ IV.2 Moreau–Yosida in inverse Kohn–Sham
While the procedure from the preceding section remains a hard task for interacting particles, density–potential inversion is often considered for the non-interacting KS system ($\lambda=0$) and then typically specified as “inverse Kohn–Sham” DFT. Starting with a given ground-state density $\rho$ for some external potential $v$ in an interacting system that is also non-interacting $v$-representable, the Kohn–Sham potential must satisfy
@@
Motivated by the fact that $\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)\to\rho$ as $\varepsilon\to 0$ one can then under technical assumptions prove that indeed $v_{\mathrm{s}}$ corresponds to the limit of $\tfrac{1}{\varepsilon}J(\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)-\rho)$ as the equations above suggest.
Now, exactly this method was proposed and implemented as a density–potential inversion scheme by Zhao, Morrison, and Parr [2] (and was later termed “ZMP method”), where the choice of spaces corresponds to homogeneous Sobolev spaces for potentials and their duals for densities (also see Section VI.1 and an upcoming publication for periodic domains like in Section VII). But in the original publication the method was developed ad hoc and only proven to work empirically. Only quite recently, it was shown [4] that the proximal-point iteration with $\varepsilon\to 0$ relates closely to the ZMP method for the usual Sobolev spaces and that it also reproduces a method by van Leeuwen and Baerends [64] if employed on a Hilbert space with $J=J^{-1}=\mathop{\mathrm{id}}$. It is further equivalent to yet another density–potential inversion method by Kumar and Harbola [65], where different choices for their penalty function $S[\rho]$ would correspond to different topologies for $X$. The method has been analyzed numerically [19; 66] (see also Section VII here) and adds itself to a long list of KS inversion techniques [67]. More than that, it shows that MY regularization offers a mathematical framework that allows for an elegant unification of numerous density–potential inversion methods.

### v2

@@ III.1 MY based on the universal density functional
The most straightforward way to introduce MY regularization into DFT follows from its definition in Eq. (15): By directly applying the method to the non-differentiable (but convex and lower semicontinuous) $F^{\lambda}(\rho)$, yielding a regularized universal density functional $F_{\varepsilon}^{\lambda}(x)$ that is well-defined and finite for all $x\in X$,
@@
where $\varepsilon>0$ is the smoothing parameter. An illustration of a regularized functional in one dimension together with the potentially non-convex $\tilde{F}^{\lambda}(x)$ and convex $F^{\lambda}(x)$ is shown in Figure 2. The regularized functional has a well-defined functional derivative (see v from the properties of MY regularization in Section II.3) and the regularization is lossless in the sense that it conserves the overall infimum of the functional (see ii).
We next introduce the corresponding energy functional $E_{\varepsilon}^{\lambda}(v)$ as the convex conjugate of $F_{\varepsilon}^{\lambda}(x)$,
@@
Note that the resulting $E_{\varepsilon}^{\lambda}(v)$ is not the MY regularization of $E^{\lambda}(v)$. Instead, the following relation holds (from Eq. (17), where the different sign is due to Eq. (32)),
@@
Since we subtract the norm-square of $v$ from the concave $E^{\lambda}(v)$, the MY energy $E_{\varepsilon}^{\lambda}(v)$ is strongly concave, which implies strict concavity. Conversely, it holds that the Legendre–Fenchel transform of a strictly concave energy functional is always differentiable.
With $v=-\nabla F_{\varepsilon}^{\lambda}(x)$ we now have a unique representing potential $v\in X^{*}$ for every $x\in X$. We here write $\nabla F_{\varepsilon}^{\lambda}$ instead of $\partial F_{\varepsilon}^{\lambda}$ because we know that the regularized functional is differentiable. This then automatically yields a well-defined and unique mapping from ‘densities’ (elements of $X$) to potentials (elements of $X^{*}$) without any $v$-representability issue. We are not even limited to the $N$-representability set $D\subset X$ any more, and Section III.3 will help to make sense of the generalized densities $x\in X$. Uniqueness of the potential also means we automatically have a Hohenberg–Kohn-type result.
In the Hilbert-space setting (Assumption 3) we further obtain a quantitative bound that shows how the difference between two potentials is bound by the difference between the corresponding densities. This comes in form of the Lipschitz estimate from property ix:
@@
This bound is lost in the limit $\varepsilon\to 0$, in agreement with the fact that, in standard DFT, two densities that lie very close to each other can come from vastly different potentials [37], making density–potential inversion ill-posed and thus notoriously difficult. A similar bound is also possible in more general Banach spaces if the MY regularization is adapted to the ‘modulus of convexity’ of the space [76]. Furthermore, MY regularization allows for a well-defined formulation of the KS iteration procedure (Section V).
[…]
@@ III.2 MY based on the total energy functional
A different viewpoint starts from the concave ground-state energy functional $E^{\lambda}(v)$ and introduces a strictly concave energy $E_{\varepsilon}^{\lambda}(v)$ by just subtracting $\varepsilon\phi(v)=\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$:
@@
Convex conjugation of this strictly concave functional then yields the regularized density functional of Eq. (49) as shown by Eq. (17),
@@
We have thus defined MY regularization starting from the energy functional instead of the universal density functional. But talking about the total energy allows to give this manipulation a physical interpretation. For a given external potential $v\in X^{*}$, we introduce a field energy functional
$E^{\mathrm{f}}(v)=-\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}},$ (55)
where the minus sign that makes the functional concave, just like the concavity of the matter energy functional $E^{\lambda}(v)$. Of course, to retrieve the correct physical electrostatic field energy, $\frac{\varepsilon_{0}}{2}\|\nabla v\|^{2}_{L^{2}}$, an appropriate choice of Banach space $X^{*}$ must be made; see Section VI.1. We then introduce the new energy functional (that could be interpreted as the total energy of matter and field)
@@
Regarding the negative sign in Eq. (55), remember that in Eq. (50) it is $F_{\varepsilon}^{\lambda}(x)+\langle v,x\rangle$ that is varied in order to find the ground state. There, the additional term is $\tfrac{1}{2\varepsilon}\|x-\rho\|^{2}_{X}$ as seen in Eq. (49) and using the relation motivated in Eq. (57) below this is equal to $\tfrac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$, precisely the expression for the (positive) field energy. Viewed in this manner, it is the energy content of the external field that regularizes $F^{\lambda}(\rho)$. Conceptually, this is an extremely attractive viewpoint, since it explains the regularization as a physical effect.
@@ III.3 MY based on density–potential mixing
Remember that we defined the $N$-representability set $D\subset X$ in Eq. (II.4) as the set of all $\rho\in X$ where $F_{\mathrm{DM}}^{\lambda}(\rho)<\infty$, known as the effective domain of $F_{\mathrm{DM}}^{\lambda}(\rho)$. For the sake of argument, we now suppose Assumption 4 under which $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$. Now even for $\rho\in D$, we can have issues with $\partial F^{\lambda}(\rho)$. First, $\partial F^{\lambda}(\rho)$ may be empty. In such cases, no potential $v\in X^{\ast}$ exists such that $v\in-\partial F^{\lambda}(\rho)\iff\rho\in\partial E^{\lambda}(v)$, implying that $\rho$ is not $v$-representable. Also, at $v$-representable densities a small variation of $\rho$ can lead outside the set of $v$-representable densities [59; 37]. There are further examples on small lattice systems where even large variations of $v$ lead to almost (or even exactly) the same ground-state density $\rho$ (and thus give counterexamples to the Hohenberg–Kohn theorem) [79; 81].
It would be better altogether if we could ignore this difficult notion of “$v$-representability” and instead are able to give a physical interpretation to all elements $x\in X$. This is achieved with the inverse duality map $J^{-1}\colon X^{*}\to X$ as the canonical way of mapping potentials to densities. For any $v$-representable $\rho\in\partial E^{\lambda}(v)$ coming from $v\in X^{*}$ define the mixed density
@@
This definition is unique only if a Hohenberg–Kohn-type result is available so that the corresponding $v\in-\partial F^{\lambda}(\rho)$ is unique, but even without uniqueness we can make this assignment and will simply allow multiple such combinations. Choosing $\varepsilon>0$ small means that in order to get the generalized density, we only mix in a small quantity of the corresponding potential $v$. We note that the resulting $x$ does not need to be positive, normalizable, or have a finite kinetic energy, only the $v$-representable density $\rho$ has all those properties. This is immediately visible from the definition in Eq. (57), where $x$ is seen to be the sum of the electronic density $\rho$ and $\varepsilon J^{-1}(v)$. With $v\in X^{*}$ as an external potential with assigned density $\varepsilon J^{-1}(v)$ that gets added to the internal electronic density of a quantum system, one could call $x$ a ‘bidensity’. It has been given different names before, such as ‘depot’ (for density–potential), ‘pseudodensity’ [54], or ‘quasidensity’ [77]. We will not commit to a singular nomenclature here, and instead refer to it as a general element $x\in X$ or a ‘mixed density’ when its decomposition like in Eq. (57) matters.
One physically motivated possibility is to choose the spaces such that $J$ corresponds to solving the Poisson equation of electrostatics and $\varepsilon J^{-1}(v)=-\varepsilon\Delta v$. Then $\rho_{\mathrm{ext}}=\varepsilon J^{-1}(v)$ can be physically interpreted as just the external density that enters the Poisson equation to yield the electrostatic potential $v$. The mixed density is then the total observable density of the system, $x=\rho-\rho_{\mathrm{ext}}$ (the minus sign is needed since we have used the convention that the electronic density is positive, $\rho\geq 0$). Interestingly, even the MY parameter $\varepsilon$ then receives a physical interpretation as the vacuum permittivity. This choice was previously suggested in the context of density–potential inversion [73]; it will be covered in more detail in Section VI.1 and has already found multiple use in ameliorating the Kohn–Sham method [Görling1999; 41; 15; 14], albeit not with the connection to the topology of density and potential space.
@@ IV Density–potential inversion
The limit procedure in property vii of the MY regularization in Section II.3 that says $\lim_{\varepsilon\to 0}\nabla F_{\varepsilon}^{\lambda}(x)\in\partial F^{\lambda}(x)$ will be the basis for a method to calculate the corresponding external potential from a given ground-state density. The method itself is already known since a long time as the “ZMP method” [112], but our derivation as well as the accompanying interpretation is only feasible by resorting to MY regularization. We also add a more speculative subsection that makes an interesting and previously unexplored connection to the Dyson equation. In this section we will require Assumption 1 but sometimes, when it is mentioned, demand higher regularity of the involved spaces for additional results. Assumption 4 ($F^{\lambda}=F_{\mathrm{DM}}^{\lambda}$) is further necessary in order to stay connected to solutions of the Schrödinger equation as explained in Section II.4.
@@
whose stationarity condition $\partial E^{\lambda}(v)+\rho\ni 0$ gives $v\in-\partial F^{\lambda}(\rho)$ as an optimizer if $\rho$ is assumed $v$-representable. This approach to density–potential inversion was first applied to atoms by Colonna and Savin [27] in 1999 and by Teale, Coriani and Helgaker to molecules [95; 96; 97] in 2009. It is sometimes called “Lieb optimization”, “Lieb maximization” [96; 97; 42] or the “dual problem” [38] in the literature. The method of Wu and Yang [109] turns this principle into a practical constrained optimization problem for the KS system. Convergence of Lieb optimization can be facilitated by making the functional strictly concave by subtracting $\frac{\varepsilon}{2}\|v\|^{2}_{X^{*}}$, which already corresponds to MY regularization,
@@
We use Eq. (58) to translate this to $\rho\in\partial E_{\varepsilon}^{\lambda}(v_{\varepsilon})$ which is equivalent to $v_{\varepsilon}=-\nabla F_{\varepsilon}^{\lambda}(\rho)$ by using the analogous relation to Eq. (39) for the regularized functionals. By property v from Section II.3, we can connect $v_{\varepsilon}$ to the proximal map,
@@ IV.1 Proximal-point iteration and relation to ZMP
As mentioned before, it is not guaranteed that this subdifferential is non-empty even for $\rho\in D\subset X$ and, in non-standard settings for DFT (like DFT on finite lattices [79]), it may even be multi-valued. For the regularized functional, no such problems occur, and we can evaluate
$v_{\varepsilon}=-\nabla F_{\varepsilon}^{\lambda}(\rho)$ (65)
@@
Due to the possibility of degeneracy in the ground state, the next step $p_{i+1}$ is possible non-unique, which creates a potential issue that has not been explored up to date. If converged, this process yields the proximal point $p=\mathop{\mathrm{prox}}_{\varepsilon F^{\lambda}}(\rho)$ and we retrieve Eq. (63),
@@
where the resulting potential $v_{\varepsilon}$ depends on the choice of the smoothing parameter $\varepsilon$. Repeating this iterative scheme for a sequence of decreasing $\varepsilon>0$, we can extrapolate to the limit $\varepsilon=0$ and obtain $v_{\varepsilon}\rightharpoonup v$. Under the minimal Assumption 1, convergence $v_{\varepsilon}\rightharpoonup v$ is weak; under the stronger Assumption 2, convergence is strong. Also by virtue of vii, the resulting potential $v$ is the element of $\partial F^{\lambda}(\rho)$ with the smallest norm.
The benefit of this method is that we do not need to evaluate the (usually unknown) $F^{\lambda}(\rho)$, relying instead on $\partial E^{\lambda}(v)$ by solving the Schrödinger equation for the ground state. This density-potential inversion procedure using MY regularization and its relation to the closely related ZMP method [112] was previously detailed in Penz et al. [73].
@@ IV.2 Moreau–Yosida in inverse Kohn–Sham
While the procedure from the preceding section remains a hard task for interacting particles, density–potential inversion is often considered for the non-interacting KS system ($\lambda=0$) and then typically specified as the “inverse Kohn–Sham method”. Starting with a given ground-state density $\rho$ for some external potential $v$ in an interacting system that is also non-interacting $v$-representable, the Kohn–Sham potential must satisfy
@@
Motivated by the fact that $\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)\to\rho$ as $\varepsilon\to 0$, one can then under technical assumptions prove that indeed $v_{\mathrm{s}}$ corresponds to the limit of $\tfrac{1}{\varepsilon}J(\mathop{\mathrm{prox}}_{\varepsilon F^{0}}(\rho)-\rho)$ as the equations above suggest.
Now exactly this method was proposed and implemented as a density–potential inversion scheme by Zhao et al. [112] (and was later termed “ZMP method”), where the choice of spaces corresponds to homogeneous Sobolev spaces for potentials and their duals for densities (also see Section VI.1 and an upcoming publication for periodic domains like in Section VII). But in the original publication the method was developed ad hoc and only proven to work empirically. Only quite recently it was shown [73] that the proximal-point iteration with $\varepsilon\to 0$ relates closely to the ZMP method for the usual Sobolev spaces and that it also reproduces a method by van Leeuwen and Baerends [102] if employed on a Hilbert space with $J=J^{-1}=\mathop{\mathrm{id}}$. It is further equivalent to yet another density–potential inversion method by Kumar and Harbola [53], where different choices for their penalty function $S[\rho]$ would correspond to different topologies for $X$. The method has been analyzed numerically [43; 11] (see also Section VII here) and adds itself to a long list of KS inversion techniques [91]. More than that, it shows that MY regularization offers a mathematical framework that allows for an elegant unification of numerous density–potential inversion methods.